Imagine a light planet revolving around a very massive star in a circular orbit of radius R with a period of revolution T. If the gravitational force of attraction between the planet and the star is proportional to R-5/2
(a) T2 is proportional to R3
(b) T2 is proportional to R7/2
(c) T2 is proportional to R3/2
(d) T2 is proportional to R3/73

Grade:6

1 Answers

Deepak Patra
9 years ago

(b) KEY CONCEPT : The centripetal force is provided by the gravitational force of attraction

mR?2 = GMmR-5/2

? mR x 4p2 / T2 = GMm /R5/2 ? T2 ? R7/2